OpenStudy (anonymous):

g (f) =square root of x+3 graph it

OpenStudy (anonymous):

If you in Real set of number, x must be >=-3. So using Matlab:

OpenStudy (anonymous):

but we do this by hand using inverted function and then rotating the graph g=sqrt(x + 3) =>> g^2 = x +3 =>> x = g^2-3, this we know how to draw! =)